# Multi-perspective Feedback-attention Coupling Model for Continuous-time Dynamic Graphs

### v1

@@ IV-A Evolving perspective
In the evolving perspective, a sefl-supervised temporal attention is designed to obtain the evolving neighborhood embedding representation of node $v_{i}$ at time t. This perspective is used to capture the real-time dynamics of the network. Given a node $v_{i}$ involved in an current event, its evoleving information at time $t$ and just before $t$ are denoted by ${h_{i}}({t})$ and ${h_{i}}({t^{-}})$, respectively. We denote by ${N_{i}}(t)=[{v_{1}},...,{v_{j}}]$ the temporal neighbors, and $[{h_{1}}(t^{-}),...,{h_{j}}(t^{-})]$ the evolving representation of neighbors just before time $t$. For the temporal self-attention, the inputs are ${h_{i}}({t^{-}})$, the current timepoint t, $[{h_{1}}(t^{-}),...,{h_{j}}(t^{-})]$, and the edge features $[{e_{i,1}}({t^{-}}),...,{e_{i,j}}({t^{-}})]$. The specific calculation process is as follows:
$\begin{split}head&=atten(Q(t),K(t),V(t))\\ &=Soft\max(\frac{{Q(t)K{{(t)}^{T}}}}{{\sqrt{d}}})V(t)\end{split}$ (2)
$P_{N_{i}(t)}^{e}=Concat(hea{d_{1}},hea{d_{2}}){W_{con}}$ (3)
The query $Q(t)={h_{i}}({t^{-}})||\Phi(0)$ is the relevant information of the target node. The key $K(t)=[{h_{1}}(t^{-})||{e_{i,1}}({t^{-}})||\Phi(t-t_{1}),\;.\;.\;.\;,{h_{j}}(t)||{e_{i,j}}({t^{-}})||\Phi(t-{t_{j}})]$ and value $V(t)=K(t)$ are the relevant information of temporal neighbors of the target node.
IV-B Raw perspective
IV-B1 Feedback coefficients with growth characteristics
For the raw perspective, we consider the factor that it is a growth process from the raw to the evolving state. After the model has grown over a period of time, the model will be more profound insights into the previous events. This insight is reflected not only in the feedback coefficients but also in the evolution information of the neighbors, which the model uses to obtain more useful information for future events from the raw interaction. In addition, During the learning process of node representation, historical interactions that contribute to the current event should not be too much disturbed by other unrelated matters. First, we compute feedback coefficients with growth properties, and then use the feedback coefficients to aggregate the raw interaction information. Given the embedding information $z_{i}(t^{-})$ and $z_{j}(t^{-})$ of nodes $v_{i}$ and $v_{j}$ just before time $t$, we compute the feedback coefficients from the time $t_{j}$ when $v_{i}$ and $v_{j}$ interacted to the present time $t$.
${I_{i,j}}(t_{j}\to t)=\psi({g_{i,j}}({t_{j}\to t})),$ (4)
where ${g_{i,j}}({t_{j}\to t})$ is used to learn the growth properties from time $t_{j}$ to $t$. Here we use only simple linear mappings, which can be computed by:
${g_{i,j}}(t_{j}\to t)=act({W_{r_{0}}}[{h_{i}}(t^{-})||{h_{j}}(t^{-})||\Phi(t-{t_{j}})]).$ (5)
Here, ||||| | is the concatenation operator, and act(.) is the active function. We stipulate that the feedback coefficient must be positive. Therefore, the external function must satisfy that its values are all greater than or equal to zero. For simplicity, we so we choose sigmoid as the external function.
${I_{i,j}}(t_{j}\to t)={(1+{e^{W_{r_{1}}{g_{i,j}}({t_{j}\to t})}})^{-1}}$ (6)
The final feedback coefficients can be obtained as follows:
@@ IV-B1 Feedback coefficients with growth characteristics
These coefficients are then utilized to aggregate the raw neighborhood information for target node i.
$P_{N_{i}(t)}^{r}=\sum\limits_{j\in{N_{i}}(t)}{{a_{i,j}}(t_{j}\to t)*{W_{trans}}\;}{C_{j}}({t_{j}}),$ (8)
where ${C_{j}}({t_{j}})=[{r_{j}}({t_{j}})||\Phi(t-{t_{j}})]$, and $W_{trans}$ is the linear transformation parameter; two linear transformations are used here. We have now finished presenting information capture from both perspectives, and the following section describes the interrelated learning of these two perspectives.
IV-C Attention coupling learning
As raw and evolving perspectives mutually drive the evolution of the temporal network, which alternately influence the learning process of network embeddings. In order to enable the model to have better performance, the evolving and raw embeddings are coupled to guide each other, which can also be regared as the interaction between the two attentions. First, we consider three types of information, i.e., changes in the node itself, raw information, and evolving information:
$z_{i}^{1}(t)={W_{node}}[{h_{i}}({t^{-}})||\Phi(t-t^{-})],$ (9)
$z_{i}^{2}(t)={W_{r}}[{h_{i}}({t^{-}})||P_{N_{i}(t)}^{e}],$ (10)
$z_{i}^{3}(t)={W_{e}}[{h_{i}}({t^{-}})||P_{N_{i}(t)}^{r}],$ (11)
Then we pass it to a two-layer feedforward neural network (FNN) for attention coupling learning.
${z_{i}}(t)=FNN([z_{i}^{1}(t)||z_{i}^{2}(t)||z_{i}^{3}(t)])$ (12)
After obtaining the embedding of nodes $v_{i}$ and $v_{j}$ at time t, the MLP decoder can be used to determine whether nodes $v_{i}$ and $v_{j}$ are interacting at time $t$, i.e., $p_{i,j}(t)=MLP(z_{i}(t),z_{j}(t))$.
@@ IV-D Training
The optimization of parameters typically depends on downstream tasks. Here link prediction task is chosen to train MPFA. Given N events arranged in chronological order, we randomly sample the same number of negative pairs from dataset that not occur in the batch. For the sake of simplicity, after a batch of events is decoded , the Binary Cross Entropy function ${l_{n}}=-({y_{n}}*\log(\delta({s_{n}})))+(1-{y_{n}})*\log(1-\delta({s_{n}})),n\in 2N$ is leveraged to train the parameters, where $y_{n}$ denotes the label of the $n^{th}$ sample pair, and the $s_{n}$ represents the probability that the $n^{th}$ sample pair is predicted to be positive. The $\delta\left({{s_{n}}}\right)$ is the sigmoid function for numerical stability.
@@ V Experiments
To validate the effectiveness of the model, we experiment extensively on a self-organizing dataset and seven public datasets. We compare it to baselines in dynamic link prediction and dynamic node classification, using both transductive and inductive inference. In addition, we perform model ablation, hyperparameter analysis, attention coefficient visualization, time consumption analysis, and long-term dependence analysis.
@@ V-A Datasets
USLegis:11footnotemark: 1 USLegis is a Senate co-sponsorship network. It records instances where members of the United States Senate co-sponsor the same bill in a given Congress. The dataset has 225 nodes and 60,396 edges.
Wikipedia:22footnotemark: 2 This dataset contains 157,474 temporal edges with 1,000 most edited pages and 8227 active users as nodes. Labels indicate whether users are prohibited from editing the page, and the interactions are transformed into text features. Both Reddit and Wikipedia are bipartite graphs.
@@
TABLE I: Results of the transductive link prediction task on MOOC, UNvote and USLegis, with AP, AUC and ACC metrics, highlighting the best-performing results for each dataset in bold. All results are shown as percentages, obtained by multiplying the values by 100 for standardization.
Dataset MOOC UNVote USLegis
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 89.37±1.6 91.17±1.3 84.13±1.5 66.13±1.2 70.31±1.9 63.97±0.9 73.91±0.7 81.37±0.7 73.71±0.6
GraphSAGE-t 84.76±1.5 87.31±1.4 80.37±1.3 65.76±0.9 70.72±1.1 63.07±0.7 71.98±0.8 78.32±0.9 71.20±0.9
JODIE 80.47±2.0 83.61±1.6 76.59±1.9 62.99±1.1 67.65±1.2 62.43±1.0 73.31±0.4 81.57±0.2 73.39±0.2
TGAT 85.09±0.4 86.20±0.4 77.40±0.4 51.22±0.6 51.49±0.9 51.07±0.7 65.75±6.0 72.55±7.8 65.16±5.1
DyRep 77.73±3.9 81.41±3.6 74.50±3.4 62.38±0.6 66.81±0.7 61.80±0.5 69.16±3.0 76.68±3.3 69.53±2.6
TGN 88.89±1.5 90.89±1.2 83.29±1.5 65.62±1.2 70.11±1.4 64.16±1.0 75.13±1.3 82.39±1.0 74.31±0.8
CAWN 69.15±0.3 70.20±0.3 63.44±0.2 51.71±0.1 52.75±0.1 50.76±0.2 69.94±0.4 75.83±0.5 66.92±0.5
EdgeBank 52.91±0.0 54.84±0.0 50.08±0.0 54.70±0.0 58.34±0.0 58.34±0.0 55.45±0.0 59.03±0.0 56.48±0.0
TCL 82.60±0.8 83.15±0.5 74.31±0.4 50.81±0.7 50.91±1.0 50.58±0.6 68.54±1.2 75.49±0.7 67.06±0.8
GraphMixer 82.00±0.2 83.31±0.2 74.56±0.1 52.28±0.5 53.31±0.5 51.92±0.3 69.28±0.9 76.09±0.6 67.40±0.7
DyGFormer 86.91±0.1 86.53±0.1 77.32±0.5 54.09±0.2 54.94±0.4 53.37±0.3 68.31±1.0 75.52±0.8 67.35±1.0
@@
TABLE II: Transductive link prediction task results on LastFM, SocialEvo and ComplexDynamics using AP, AUC and ACC metrics, with the best results for each dataset highlighted in bold. All results are presented as percentages, obtained by multiplying the values by 100 for standardization.
Dataset LastFM SocialEvo ComplexDynamics
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 70.37±1.7 71.29±1.3 67.46±1.3 90.33±0.2 91.36±0.9 87.41±1.5 88.12±0.1 87.04±0.1 78.54±0.2
GraphSAGE-t 73.21±1.8 74.32±1.3 67.76±0.9 90.21±1.5 92.59±0.9 86.29±1.3 86.95±0.2 85.67±0.3 77.34±0.4
JODIE 70.21±1.9 69.88±1.4 64.83±0.9 88.55±1.8 90.96±1.3 83.77±1.5 82.93±0.8 82.67±0.6 74.59±0.6
TGAT 72.43±0.3 70.69±0.2 64.52±0.2 92.71±0.7 94.32±0.5 89.32±0.3 84.57±0.2 83.01±0.1 74.49±0.2
DyRep 70.55±2.1 70.17±1.9 65.00±1.3 85.11±0.5 89.05±0.3 82.13±0.2 79.17±0.7 78.77±0.4 71.66±0.3
TGN 75.29±3.7 76.41±3.2 69.91±2.6 93.70±0.3 95.68±0.2 90.32±0.4 87.84±0.3 86.82±0.2 78.35±0.3
CAWN 73.36±0.2 69.43±0.1 63.48±0.2 78.46±0.3 80.43±0.4 74.89±0.6 90.25±0.1 88.35±0.2 79.39±0.2
EdgeBank 77.25±0.0 83.51±0.0 82.06±0.0 52.04±0.0 53.90±0.0 53.90±0.0 53.90±0.0 70.82±0.0 50.00±0.0
TCL 69.39±2.6 64.67±2.1 59.47±1.7 93.07±0.2 94.38±0.2 90.22±0.1 89.18±0.1 87.30±0.1 78.32±0.1
GraphMixer 75.68±0.2 73.42±0.3 67.21±0.2 92.56±0.4 94.47±0.3 89.35±0.2 87.04±0.2 85.79±0.2 77.50±0.1
DyGFormer 83.81±0.2 79.97±0.2 71.63±0.3 94.21±0.1 95.56±0.2 91.83±0.2 89.28±0.1 86.84±0.1 77.55±0.1
@@
TABLE III: Results of the inductive link prediction task on MOOC, UNvote and USLegis, with AP, AUC and ACC metrics, highlighting the best-performing results for each dataset in bold. All results are shown as percentages, obtained by multiplying the values by 100 for standardization.
Dataset MOOC UNVote USLegis
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 87.36±2.5 88.76±2.4 81.32±2.7 57.71±3.4 59.21±4.6 55.73±3.6 59.16±1.7 60.93±1.3 57.78±0.7
GraphSAGE-t 84.23±1.5 86.72±1.2 78.32±1.1 57.65±1.3 58.97±1.4 55.37±1.9 60.31±0.7 59.37±0.4 57.31±0.6
JODIE 80.54±0.9 83.44±0.9 76.43±0.9 55.74±1.7 56.65±2.7 53.58±2.12 52.16±0.5 55.07±0.8 53.15±0.5
TGAT 84.53±0.5 85.68±0.4 77.02±0.5 52.21±1.0 51.96±1.1 51.61±0.76 49.71±2.3 47.33±3.5 48.68±2.9
DyRep 78.52±2.8 82.11±2.3 74.99±2.2 54.38±1.6 54.77±2.3 52.37±1.81 56.26±2.0 58.87±2.1 55.56±1.6
TGN 88.79±1.4 90.82±1.3 83.25±1.5 57.14±2.6 58.90±4.0 55.28±3.00 59.52±1.3 62.14±1.7 58.26±1.0
CAWN 68.27±0.6 68.66±0.7 61.91±0.6 49.74±0.7 48.32±0.6 47.67±0.33 53.11±0.4 51.01±0.5 50.11±0.8
TCL 80.95±0.8 81.57±0.5 72.98±0.4 50.20±1.0 49.90±1.5 49.32±1.16 49.43±1.0 46.84±1.1 48.19±1.0
GraphMixer 80.56±0.2 81.99±0.1 73.22±0.2 51.01±0.8 50.97±0.9 50.17±0.7 49.31±1.5 46.02±1.3 47.23±1.3
DyGFormer 85.66±0.2 85.20±0.2 76.05±0.2 53.87±0.2 54.31±0.2 53.16±0.24 48.08±0.6 45.34±0.9 46.84±1.0
@@
TABLE IV: Inductive link prediction task results on LastFM, SocialEvo and ComplexDynamics using AP, AUC and ACC metrics, with the best results for each dataset highlighted in bold. All results are presented as percentages, obtained by multiplying the values by 100 for standardization.
Dataset LastFM SocialEvo ComplexDynamics
Metric AP AUC ACC AP AUC ACC AP AUC ACC
GAT-t 79.73±2.1 80.32±1.7 73.18±1.1 82.33±0.2 83.21±3.1 76.36±3.1 80.15±0.3 78.49±0.3 70.75±0.4
GraphSAGE-t 81.32±1.3 81.29±0.7 75.63±0.5 79.66±5.7 83.57±4.5 77.23±4.0 80.26±0.7 78.37±0.8 70.73±0.6
JODIE 82.36±1.1 81.63±1.0 75.34±0.8 91.22±1.9 93.10±1.4 87.68±1.8 73.90±1.0 71.63±1.0 65.44±0.7
TGAT 77.74±0.4 76.27±0.2 69.90±0.8 91.25±0.4 93.10±0.8 87.54±1.0 83.00±0.3 80.87±0.3 72.76±0.4
DyRep 81.65±1.8 81.01±1.9 74.75±1.4 82.37±0.6 84.87±0.3 78.56±0.2 67.80±1.1 66.77±1.0 61.73±0.7
TGN 80.58±3.8 81.51±3.0 75.26±2.6 90.70±0.7 93.37±0.4 89.32±0.5 79.40±0.6 77.53±0.6 69.57±0.4
CAWN 77.36±0.3 73.16±0.4 67.43±0.3 77.17±0.3 79.46±0.4 74.67±1.0 87.37±0.1 84.28±0.2 75.10±0.4
TCL 72.07±3.2 68.76±2.9 63.21±2.7 91.97±0.3 93.13±0.2 89.53±0.2 87.86±0.1 85.31±0.1 76.53±0.3
GraphMixer 82.14±0.4 80.23±0.3 73.99±0.2 91.23±0.4 93.28±0.3 88.13±0.2 80.95±0.4 78.24±0.4 70.84±0.3
DyGFormer 86.96±0.2 83.82±0.3 76.17±0.3 92.26±0.1 94.32±0.2 90.67±0.2 88.25±0.1 85.13±0.1 76.01±0.3
@@
TABLE V: Dynamic node classification with AUC. All results have been converted to percentages by multiplying by 100.
Wikipedia Reddit
GAT 82.34±0.8 64.52±0.5
GraphSAGE 82.42±0.7 61.24±0.6
CTDNE 75.89±0.5 59.43±0.6
JODIE 84.84±1.2 61.83±2.7
TGAT 83.69±0.7 65.56±0.7
DyRep 84.59±2.2 62.91±2.4
TGN 87.81±0.3 67.06±0.9
CAWN 84.88±1.3 66.34±1.8
TCL 77.83±2.1 68.87±2.2
GraphMixer 86.80±0.8 64.22±3.3
DyGFormer 87.44±1.1 68.00±1.7

### v2

@@ IV-A Evolving perspective
In the evolving perspective, we employ a multi-head temporal self-attention module to obtain the fusion of the current state of historical interactions of node $v_{i}$ at time $t$, aiming to capture the real-time dynamics of the network. Specifically, the current states of node $v_{i}$ and its historical interaction nodes at time $t$ are represented as $h_{i}\left(t\right)$ and $\left[h_{1}\left(t\right),\ldots,h_{j}\left(t\right)\right]$, respectively. Then, the fusion of the current states of historical interaction events of node $v_{i}$ at time $t$, denoted as $P_{N_{i}\left(t\right)}^{e}$, is calculated as follows,
$P_{N_{i}(t)}^{e}=$ $MultiHeadAtten(Q(t),K(t),V(t))$ (4)
$=Softmax{(\frac{Q(t)K(t)^{T}}{\sqrt{d}})}V(t),$
where $N_{i}\left(t\right)$ denotes the set of historical interaction events of node $v_{i}$ at time $t$ (i.e., the temporal neighbors). $Q(t)$, $K(t)$ and $V(t)$ are represented as follows,
$Q(t)=h_{i}(t)||\Phi(0)$ (5)
$K(t)=$ $[h_{1}(t)||e_{i,1}(t)||\Phi(t-t_{1}),$ (6)
$...,h_{j}(t)||e_{i,j}(t)||\Phi(t-t_{j})],$
$V\left(t\right)=K\left(t\right).$ (7)
IV-B Original perspective
The original perspective considers the essence of node interactions. To obtain the fusion of the original states of the node’s historical interaction events, we first compute a feedback attention coefficient with growth characteristics. This coefficient can be viewed as the model’s growth process from the original state to the current state. After a period of model growth, the model gains deeper insights into previous events. Specifically, given a historical interaction of node $v_{i}$ at time $t_{j}$ represented by $\left(v_{i},v_{j},e_{i,j}\left(t_{j}\right)\right)$, we compute the feedback attention coefficient from time $t_{j}$ to the current time $t$,
$I_{i,j}(t_{j}\rightarrow t)=\psi(g_{i,j}(t_{j}\rightarrow t)),$ (8)
here, $g_{i,j}(t_{j}\rightarrow t)$ is utilized to learn the growth characteristics from time $t_{j}$ to $t$. A single layer of non-linear mapping is employed, computed as follows,
${g_{i,j}}(t_{j}\to t)=relu({W_{1}}[{h_{i}}(t)||{h_{j}}(t)||\Phi(t-{t_{j}})]),$ (9)
where $w_{1}$ is a trainable parameter, ||||| | is the concatenation operator, and $relu()$ is the activation function. Since the feedback coefficients must be between 0 and 1, the external function $\psi$ must satisfy that all of its values are greater than or equal to zero. Therefore, the sigmoid function is chosen as the external function, i.e,
${I_{i,j}}(t_{j}\to t)={(1+{e^{W_{2}{g_{i,j}}({t_{j}\to t})}})^{-1}},$ (10)
where $w_{2}$ is a trainable parameter. The final feedback coefficient can be obtained as follows,
@@ IV-B Original perspective
Next, we aggregate the original state information of the historical interactions of the target node $v_{i}$ using the obtained feedback coefficients,
$P_{N_{i}(t)}^{r}=\sum\limits_{j\in{N_{i}}(t)}{{a_{i,j}}(t_{j}\to t)*{W_{trans}}\;}[{r_{j}}({t_{j}})||\Phi(t-{t_{j}})],$ (12)
where $W_{trans}$ represents the parameters for the linear transformation. So far, we have managed to learn the evolution process of dynamic graphs from two perspectives. Next, we will introduce the mutual coupling of these two perspectives.
IV-C Multi-perspective attention coupling learning
To improve the robustness and generalization ability of the model, we perform coupling learning on the information obtained from the two perspectives to obtain the final embedding representation of the nodes. Specifically, we further parameterize the changes in the node itself, the original state, and the current state, and feed them into a two-layer feedforward neural network (FNN) to obtain the final embedding representation $z_{i}\left(t\right)$ of the node.
$z_{i}^{n}(t)={W_{n}}[{h_{i}}({t})||\Phi(t-t^{-})],$ (13)
$z_{i}^{r}(t)={W_{r}}[{h_{i}}({t})||P_{N_{i}(t)}^{e}],$ (14)
$z_{i}^{e}(t)={W_{e}}[{h_{i}}({t})||P_{N_{i}(t)}^{r}],$ (15)
$z_{i}\left(t\right)=FNN\left(\left[z_{i}^{n}\left(t\right)||z_{i}^{r}\left(t\right)||z_{i}^{e}\left(t\right)\right]\right),$ (16)
where $W_{n}$, $W_{r}$ and $W_{e}$ denote trainable parameters.
@@ IV-D Training
We employ link prediction loss to train the proposed model. After obtaining the embedding representations $z_{i}(t)$ and $z_{j}(t)$ for nodes $v_{i}$ and $v_{j}$ at time $t$, we feed them into an MLP decoder to determine whether nodes $v_{i}$ and $v_{j}$ interact at time $t$, i.e., $p_{i,j}\left(t\right)=MLP\left(z_{i}\left(t\right),z_{j}\left(t\right)\right)$. Link prediction is inherently a binary classification problem, thus we directly utilize binary cross-entropy function as the link prediction loss function, ${l_{n}}=-({y_{n}}*\log(\delta({s_{n}})))+(1-{y_{n}})*\log(1-\delta({s_{n}})),n\in 2N$, where $y_{n}$ denotes the label of the $n^{th}$ sample pair, and the $s_{n}$ represents the probability that the $n^{th}$ sample pair is predicted to be positive. The $\delta\left({{s_{n}}}\right)$ is the sigmoid function for numerical stability.
@@ V Experiments
To verify the effectiveness of our proposed model, this section conducts extensive experiments on one self-organized datasets and seven public datasets. We demonstrate the advantages of MPFA over existing algorithms on dynamic link prediction and dynamic node classification tasks. Additionally, we conduct an in-depth analysis of each component of MPFA and its intrinsic characteristics, including model ablation research, hyperparameter analysis, attention coefficient visualization, time efficiency analysis, and long-term dependency analysis.
@@ V-A Datasets
USLegis:11footnotemark: 1 USLegis is a Senate co-sponsorship network. It records instances where members of the United States Senate co-sponsored the same bill in a given Congress. The dataset has 225 nodes and 60,396 edges.
Wikipedia:22footnotemark: 2 This dataset contains 157,474 temporal edges with 1,000 most edited pages and 8227 active users as nodes. Labels indicate whether users are banned from editing the page, and interactions are transformed into text features. Both Reddit and Wikipedia are bipartite graphs.
@@
TABLE I: Results of the transductive link prediction task on MOOC, UNVOTE and USLEGIS, with AP(%), AUC(%) and ACC(%) metrics, highlighting the best-performing results for each dataset in bold. All results are shown as percentages, obtained by multiplying the values by 100 for standardization.
MOOC UNVote USLegis
Model AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%)
GAT[8]-t 89.37±1.6 91.17±1.3 84.13±1.5 66.13±1.2 70.31±1.9 63.97±0.9 73.91±0.7 81.37±0.7 73.71±0.6
GraphSAGE[9]-t 84.76±1.5 87.31±1.4 80.37±1.3 65.76±0.9 70.72±1.1 63.07±0.7 71.98±0.8 78.32±0.9 71.20±0.9
JODIE[24] 80.47±2.0 83.61±1.6 76.59±1.9 62.99±1.1 67.65±1.2 62.43±1.0 73.31±0.4 81.57±0.2 73.39±0.2
TGAT[26] 85.09±0.4 86.20±0.4 77.40±0.4 51.22±0.6 51.49±0.9 51.07±0.7 65.75±6.0 72.55±7.8 65.16±5.1
DyRep[21] 77.73±3.9 81.41±3.6 74.50±3.4 62.38±0.6 66.81±0.7 61.80±0.5 69.16±3.0 76.68±3.3 69.53±2.6
TGN[25] 88.89±1.5 90.89±1.2 83.29±1.5 65.62±1.2 70.11±1.4 64.16±1.0 75.13±1.3 82.39±1.0 74.31±0.8
CAWN[29] 69.15±0.3 70.20±0.3 63.44±0.2 51.71±0.1 52.75±0.1 50.76±0.2 69.94±0.4 75.83±0.5 66.92±0.5
EdgeBank[30] 52.91±0.0 54.84±0.0 50.08±0.0 54.70±0.0 58.34±0.0 58.34±0.0 55.45±0.0 59.03±0.0 56.48±0.0
TCL[27] 82.60±0.8 83.15±0.5 74.31±0.4 50.81±0.7 50.91±1.0 50.58±0.6 68.54±1.2 75.49±0.7 67.06±0.8
GraphMixer[39] 82.00±0.2 83.31±0.2 74.56±0.1 52.28±0.5 53.31±0.5 51.92±0.3 69.28±0.9 76.09±0.6 67.40±0.7
DyGFormer[28] 86.91±0.1 86.53±0.1 77.32±0.5 54.09±0.2 54.94±0.4 53.37±0.3 68.31±1.0 75.52±0.8 67.35±1.0
@@
TABLE II: Transductive link prediction task results on LASTFM, SOCIALEVO and COMPLEXDYNAMICS.
LastFM SocialEvo ComplexDynamics
Model AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%)
GAT[8]-t 70.37±1.7 71.29±1.3 67.46±1.3 90.33±0.2 91.36±0.9 87.41±1.5 88.12±0.1 87.04±0.1 78.54±0.2
GraphSAGE[9]-t 73.21±1.8 74.32±1.3 67.76±0.9 90.21±1.5 92.59±0.9 86.29±1.3 86.95±0.2 85.67±0.3 77.34±0.4
JODIE[24] 70.21±1.9 69.88±1.4 64.83±0.9 88.55±1.8 90.96±1.3 83.77±1.5 82.93±0.8 82.67±0.6 74.59±0.6
TGAT[26] 72.43±0.3 70.69±0.2 64.52±0.2 92.71±0.7 94.32±0.5 89.32±0.3 84.57±0.2 83.01±0.1 74.49±0.2
DyRep[21] 70.55±2.1 70.17±1.9 65.00±1.3 85.11±0.5 89.05±0.3 82.13±0.2 79.17±0.7 78.77±0.4 71.66±0.3
TGN[25] 75.29±3.7 76.41±3.2 69.91±2.6 93.70±0.3 95.68±0.2 90.32±0.4 87.84±0.3 86.82±0.2 78.35±0.3
CAWN[29] 73.36±0.2 69.43±0.1 63.48±0.2 78.46±0.3 80.43±0.4 74.89±0.6 90.25±0.1 88.35±0.2 79.39±0.2
EdgeBank[30] 77.25±0.0 83.51±0.0 82.06±0.0 52.04±0.0 53.90±0.0 53.90±0.0 53.90±0.0 70.82±0.0 50.00±0.0
TCL[27] 69.39±2.6 64.67±2.1 59.47±1.7 93.07±0.2 94.38±0.2 90.22±0.1 89.18±0.1 87.30±0.1 78.32±0.1
GraphMixer[39] 75.68±0.2 73.42±0.3 67.21±0.2 92.56±0.4 94.47±0.3 89.35±0.2 87.04±0.2 85.79±0.2 77.50±0.1
DyGFormer[28] 83.81±0.2 79.97±0.2 71.63±0.3 94.21±0.1 95.56±0.2 91.83±0.2 89.28±0.1 86.84±0.1 77.55±0.1
@@
TABLE III: Results of the inductive link prediction task on MOOC, UNVOTE and USLEGIS.
MOOC UNVote USLegis
Model AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%)
GAT[8]-t 87.36±2.5 88.76±2.4 81.32±2.7 57.71±3.4 59.21±4.6 55.73±3.6 59.16±1.7 60.93±1.3 57.78±0.7
GraphSAGE[9]-t 84.23±1.5 86.72±1.2 78.32±1.1 57.65±1.3 58.97±1.4 55.37±1.9 60.31±0.7 59.37±0.4 57.31±0.6
JODIE[24] 80.54±0.9 83.44±0.9 76.43±0.9 55.74±1.7 56.65±2.7 53.58±2.12 52.16±0.5 55.07±0.8 53.15±0.5
TGAT[26] 84.53±0.5 85.68±0.4 77.02±0.5 52.21±1.0 51.96±1.1 51.61±0.76 49.71±2.3 47.33±3.5 48.68±2.9
DyRep[21] 78.52±2.8 82.11±2.3 74.99±2.2 54.38±1.6 54.77±2.3 52.37±1.81 56.26±2.0 58.87±2.1 55.56±1.6
TGN[25] 88.79±1.4 90.82±1.3 83.25±1.5 57.14±2.6 58.90±4.0 55.28±3.00 59.52±1.3 62.14±1.7 58.26±1.0
CAWN[29] 68.27±0.6 68.66±0.7 61.91±0.6 49.74±0.7 48.32±0.6 47.67±0.33 53.11±0.4 51.01±0.5 50.11±0.8
TCL[27] 80.95±0.8 81.57±0.5 72.98±0.4 50.20±1.0 49.90±1.5 49.32±1.16 49.43±1.0 46.84±1.1 48.19±1.0
GraphMixer[39] 80.56±0.2 81.99±0.1 73.22±0.2 51.01±0.8 50.97±0.9 50.17±0.7 49.31±1.5 46.02±1.3 47.23±1.3
DyGFormer[28] 85.66±0.2 85.20±0.2 76.05±0.2 53.87±0.2 54.31±0.2 53.16±0.24 48.08±0.6 45.34±0.9 46.84±1.0
@@
TABLE IV: Inductive link prediction task results on LASTFM, SOCIALEVO and COMPLEXDYNAMICS.
LastFM SocialEvo ComplexDynamics
Model AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%) AP(%) AUC(%) ACC(%)
GAT[8]-t 79.73±2.1 80.32±1.7 73.18±1.1 82.33±0.2 83.21±3.1 76.36±3.1 80.15±0.3 78.49±0.3 70.75±0.4
GraphSAGE[9]-t 81.32±1.3 81.29±0.7 75.63±0.5 79.66±5.7 83.57±4.5 77.23±4.0 80.26±0.7 78.37±0.8 70.73±0.6
JODIE[24] 82.36±1.1 81.63±1.0 75.34±0.8 91.22±1.9 93.10±1.4 87.68±1.8 73.90±1.0 71.63±1.0 65.44±0.7
TGAT[26] 77.74±0.4 76.27±0.2 69.90±0.8 91.25±0.4 93.10±0.8 87.54±1.0 83.00±0.3 80.87±0.3 72.76±0.4
DyRep[21] 81.65±1.8 81.01±1.9 74.75±1.4 82.37±0.6 84.87±0.3 78.56±0.2 67.80±1.1 66.77±1.0 61.73±0.7
TGN[25] 80.58±3.8 81.51±3.0 75.26±2.6 90.70±0.7 93.37±0.4 89.32±0.5 79.40±0.6 77.53±0.6 69.57±0.4
CAWN[29] 77.36±0.3 73.16±0.4 67.43±0.3 77.17±0.3 79.46±0.4 74.67±1.0 87.37±0.1 84.28±0.2 75.10±0.4
TCL[27] 72.07±3.2 68.76±2.9 63.21±2.7 91.97±0.3 93.13±0.2 89.53±0.2 87.86±0.1 85.31±0.1 76.53±0.3
GraphMixer[39] 82.14±0.4 80.23±0.3 73.99±0.2 91.23±0.4 93.28±0.3 88.13±0.2 80.95±0.4 78.24±0.4 70.84±0.3
DyGFormer[28] 86.96±0.2 83.82±0.3 76.17±0.3 92.26±0.1 94.32±0.2 90.67±0.2 88.25±0.1 85.13±0.1 76.01±0.3
@@
TABLE V: Dynamic node classification with AUC(%). All results have been converted to percentages by multiplying by 100.
      Model       Wikipedia       Reddit
      GAT[8]-t       82.95±0.7       64.76±0.6
      GraphSAGE[9]-t       82.87±0.6       61.31±0.7
      JODIE[24]       84.84±1.2       61.83±2.7
      TGAT[26]       83.69±0.7       65.56±0.7
      DyRep[21]       84.59±2.2       62.91±2.4
      TGN[25]       87.81±0.3       67.06±0.9
      CAWN[29]       84.88±1.3       66.34±1.8
      TCL[27]       77.83±2.1       68.87±2.2
      GraphMixer[39]       86.80±0.8       64.22±3.3
      DyGFormer[28]       87.44±1.1       68.00±1.7